# Secure numerical simulations using fully homomorphic encryption

### v2

@@ 7 Conclusions and outlook
In this work, we presented the first prototype of secure numerical simulations using fully homomorphic encryption. We gave an overview of the CKKS scheme and its efficient implementation within the OpenFHE library, and introduced our OpenFHE.jl and SecureArithmetic.jl packages, which provide a user-friendly interface to OpenFHE in the Julia programming language. A detailed analysis of the accuracy and performance of the individual CKKS routines revealed that while each operation introduces some numerical error, bootstrapping is by far the least accurate procedure. Similar observations were made for the runtime behavior, where the bootstrapping mechanism is orders of magnitude slower than other CKKS operations.
@@
We then conducted homomorphically encrypted numerical simulations by building upon the tools and methods developed in the previous sections. Our results illustrate that it is feasible to implement secure numerical simulations with the CKKS scheme, with results that match the unencrypted simulations well. However, the encrypted operations introduce a large computational overhead and are a significant source of error, to the point that we were able to achieve less than single-precision floating point accuracy at the cost of a six-orders-of-magnitude increase in runtime. Nearly all of these undesirable effects can be attributed to the bootstrapping procedure, which is the most inaccurate and slowest operation in the CKKS scheme. However, bootstrapping is also necessary to refresh the multiplicative depth of the ciphertexts, making it indispensable for long-running simulations. Moreover, iterative bootstrapping can further improve the accuracy of an FHE algorithm, though at a non-negligible runtime penalty that needs to be taken into account. We also showed that the CKKS scheme implementation in OpenFHE is parallelizable, and that using multiple threads can lead to a significant reduction in the solution time.
@@
To summarize, our work demonstrates that it is possible to perform secure numerical simulations using fully homomorphic encryption. At the same time, the large computational overhead and precision loss in the bootstrapping operation are significant obstacles that need to be taken into account. Therefore, currently FHE-secured numerical simulations only make practical sense in specialized applications where the security of the data is of utmost importance. Thus, more detailed investigations are necessary to optimize the use of the CKKS scheme for numerical simulations. By providing a user-friendly interface to the OpenFHE library in Julia, and by implementing some important building blocks such as the $\operatorname{circshift}$ algorithms, we hope to lay the foundation for further research in this direction. In addition, our discussion of the challenges and potential solutions for extending the secure numerical simulations to other mathematical models and numerical methods should provide a starting point for other researchers interested in this field.
In the future, we plan to extend our work in several areas. We would like to increase the complexity of the simulations by looking at systems of equations and at three-dimensional problems. This will also necessitate an extension of the $\operatorname{circshift}$ operation to support multiple ciphertexts and tensors. Furthermore, we will investigate the discretization of nonlinear equations, which will require us to handle additional arithmetic operations such as division and exponentiation. Moreover, we would like to better understand the observed behavior that the accuracy of secure numerical simulations with bootstrapping can improve over time. Finally, we plan to continue developing the SecureArithmetic.jl package as a tool for experimentation and rapid prototyping with encrypted algorithms, to open up FHE to a wider audience in the computational science community.

### v3

@@ 6.4 General strategies for improving performance
6.5 Unsupported simulation types
The presented prototypes demonstrate the viability of FHE-secured numerical simulations for some simple schemes, and the previous sections outline challenges and solutions for extending them to more complex use cases. Some types of numerical simulations, however, remain infeasible—either due to inherent limitations of FHE/CKKS or due to current technical constraints.
Fundamentally unsupported are numerical methods that rely on data-dependent control flow, such as adaptivity (adaptive time stepping, adaptive mesh refinement), nonlinear solvers with conditional logic, event-driven simulations, or iterative schemes with convergence-based termination. Similarly, algorithms requiring irregular data structures or random memory access patterns are difficult or impossible to realize with the CKKS scheme. Examples include irregular sparse linear algebra, mesh-free particle methods, or multilevel algorithms with non-uniform transfer operators. Finally, numerical methods that require high numerical accuracy and are sensitive to rounding errors (typically double precision or higher), such as time integration for stiff PDEs, spectral methods, or eigenvalue solvers for ill-conditioned systems, are not well-suited for use with CKKS if they require bootstrapping.
While some of these methods may be rewritten to avoid the limitations imposed by FHE through careful algorithm design, this usually introduces an even higher computational overhead. Thus, they represent significant challenges for many common numerical methods.
@@ 7 Conclusions and outlook
In this work, we present the first prototype of secure numerical simulations using fully homomorphic encryption. We gave an overview of the CKKS scheme and its efficient implementation within the OpenFHE library, and introduced our OpenFHE.jl and SecureArithmetic.jl packages, which provide a user-friendly interface to OpenFHE in the Julia programming language. A detailed analysis of the accuracy and performance of the individual CKKS routines revealed that while each operation introduces some numerical error, bootstrapping is by far the least accurate procedure. Similar observations were made for the runtime behavior, where the bootstrapping mechanism is orders of magnitude slower than other CKKS operations.
@@
We then conducted homomorphically encrypted numerical simulations by building upon the tools and methods developed in the previous sections. Our results illustrate that it is feasible to implement secure numerical simulations with the CKKS scheme, with results that match the unencrypted simulations well. However, the encrypted operations introduce a large computational overhead and are a significant source of error, to the point that we were able to achieve less than single-precision floating point accuracy at the cost of a six-orders-of-magnitude increase in runtime. Nearly all of these undesirable effects can be attributed to the bootstrapping procedure, which is the most inaccurate and slowest operation in the CKKS scheme. However, bootstrapping is necessary to refresh the multiplicative depth of the ciphertexts, making it indispensable for long-running simulations. Moreover, iterative bootstrapping further improves the accuracy of an FHE algorithm, though at a non-negligible runtime penalty that needs to be taken into account. We also showed that the CKKS scheme implementation in OpenFHE is parallelizable, and that using multiple threads can lead to a significant reduction in the solution time.
@@
To summarize, our work demonstrates the feasibility of secure numerical simulations using fully homomorphic encryption. We developed a user-friendly interface to the OpenFHE library in Julia, introduced building blocks such as the $\operatorname{circshift}$ algorithms, and verified that the overall simulation results agree well with unencrypted baselines. These contributions lay the foundation for further research in FHE-secured scientific computing.
Looking ahead, several critical research challenges must be overcome to make secure simulations practical at scale. First, the numerical accuracy achievable with CKKS is inherently limited. Higher precision, e.g., through iterative bootstrapping, demands significantly more computational resources, and double-precision accuracy appears out of reach given the current bootstrapping implementation. Second, the homomorphic paradigm precludes core algorithmic structures in scientific computing, such as data-dependent control flow (necessary for iterative or adaptive schemes) or irregular data structures (e.g., sparse matrices). Third, the runtime overhead remains substantial—not only compared to unencrypted simulations but also in absolute terms. Bridging this performance gap will require improvements in CKKS implementations, new approaches like functional bootstrapping, and the efficient use of accelerators such as GPUs [68, 69, 70]. Emerging hardware architectures tailored for FHE workloads, including FPGAs [71, 72] and FHE-specific ASICs [73, 74], could play a central role in reducing runtimes. Addressing these limitations is the key to enabling broader use of FHE in scientific computing. In the meantime, FHE-secured numerical simulations only make practical sense in specialized applications where data confidentiality is of utmost importance.
Despite these open challenges, we believe that progress can be made incrementally. In particular, we plan to extend our work in several areas. We would like to increase the complexity of the simulations by looking at systems of equations and at three-dimensional problems. This also necessitates an extension of the $\operatorname{circshift}$ operation to support multiple ciphertexts and tensors. Furthermore, we will investigate the discretization of nonlinear equations, which requires us to handle additional arithmetic operations such as division and exponentiation. Moreover, we would like to better understand the observed behavior that the accuracy of secure numerical simulations with bootstrapping can improve over time. Finally, we plan to continue developing the SecureArithmetic.jl package as a tool for experimentation and rapid prototyping with encrypted algorithms, to open up FHE to a wider audience in the computational science community.
@@ Acknowledgments
Appendix A OpenFHE implementation details
In the OpenFHE library, the value returned by the GetLevel() function does not represent the ciphertext level as it is commonly used in literature and in this paper. Instead, it starts at zero for a fresh ciphertext and increases by one with each multiplication. Also, when using the OpenFHE library with the FLEXIBLEAUTO scaling technique as we are here (see Table 1 above), the rescaling step after a ciphertext multiplication is not performed immediately but only just before the next multiplication. Therefore, the ciphertext level reported by GetLevel() function may be off by one. These implementation-specific peculiarities must be taken into account when implementing algorithms with OpenFHE.